# Panoramic Direct LiDAR-assisted Visual Odometry

### v1

@@ Abstract
Enhancing visual odometry by exploiting sparse depth measurements from LiDAR is a promising solution for improving tracking accuracy of an odometry. Most existing works utilize a monocular pinhole camera, yet could suffer from poor robustness due to less available information from limited field-of-view (FOV). This paper proposes a panoramic direct LiDAR-assisted visual odometry, which fully associates the 360-degree FOV LiDAR points with the 360-degree FOV panoramic image datas. 360-degree FOV panoramic images can provide more available information, which can compensate inaccurate pose estimation caused by insufficient texture or motion blur from a single view. In addition to constraints between a specific view at different times, constraints can also be built between different views at the same moment. Experimental results on public datasets demonstrate the benefit of large FOV of our panoramic direct LiDAR-assisted visual odometry to state-of-the-art approaches.
@@ I Introduction
Using 3D light detection and ranging (LiDAR) information to assist visual odometry (VO), i.e., LiDAR-assisted VO, has attracted increasing interests recently due to the ability of achieving accurate pose estimation in outdoor environments. Several notable LiDAR-assisted VO systems [1, 2, 3, 4, 5, 6] fuse measurements from a monocular pinhole camera and a 360-degree field-of-view (FOV) LiDAR to perform pose estimation. These approaches suffer from poor robustness if the single view appears motion blur or insufficient texture. Compared to the monocular camera, the panoramic camera provides a 360-degree FOV image data (as shown in Fig. 1), which can more fully associate with LiDAR points and hold great potential to enhance the robustness of LiDAR-assisted VO. Compared to utilizing rigidly coupled multi-camera systems to achieve 360-degree FOV, the panoramic camera has the advantage of not requiring online extrinsic calibration and temporal synchronization between multiple cameras.
Despite a panoramic camera can provide 360-degree FOV through multiple synchronized surround-view cameras, each camera has a low frame rate and a very limited horizontal FOV. When the camera moves quickly, there is little overlap between consecutive frames from the same perspective. This limited overlap results in less information available for constructing constraints, which in turn leads to poor accuracy in pose estimation.
@@ II Related Work
Wide FOV camera-based VO. Zhang et. al. [9] conducted an in-depth investigation into the impact of camera FOV on the performance of VO, and pointed out that it is advantageous to use a large FoV camera (e.g., fisheye or catadioptric) for indoor scenes. Caruso et. al. [10] proposed a direct monocular SLAM method for omnidirectional or wide FOV fisheye cameras, which allow to observe and reconstruct a larger portion of the surrounding environment, and also make the system more robust to degenerate (rotation-only) movement. Matsuki et. al. [11] proposed a direct monocular visual odometry for omnidirectional cameras. [11] utilized the unified omnidirectional model as a projection function, which can be applied to fisheye cameras with a FoV well above 180 degrees. CubemapSLAM [12] presented a real-time feature-based SLAM system for fisheye cameras featured by a large FoV, which increase visual overlap between consecutive frames and capture more pixels belonging to the static parts of the environment. PALVO [13] applied panoramic annular lens to visual odometry, greatly increasing the robustness to rapid motion and dynamic scenarios.
@@
LiDAR-assisted VO. DEMO [1] proposed a feature-based LiDAR-assisted VO, which associates extracted 2D features with 3D points in order to assign depth value to each feature point (i.e., 2D-3D data association). LIMO [2] adds a loop closure module on the base of DEMO and utilizes semantic information to identify moving objects and reject outliers. [3] proposes to use line features in addition to point features for tracking and mapping, so as to improve the robustness to noises [24], large viewpoint changes [25], and motion blurs [26]. DVL-SLAM [4] proposed to use the direct method for LiDAR-assisted visual tracking, aiming to avoid interpolation error in 2D-3D data association.
@@ III-A Coordinate Systems
There are several types of panoramic cameras, and we specifically refer to the type that contains five surround-view cameras in this work (i.e., ). We denote , , , as a 3D point in the camera coordinate, the LiDAR coordinate, the body coordinate and the world coordinate respectively. The body coordinate is coinciding with , and is the transformation form to . The world coordinate system coincides with the body coordinate system at the initial time.
@@ III-B Camera Projection Model
where represents the depth of 2D point in the camera coordinates.
@@ VI Experiments
In this section, we evaluate our panoramic direct LiDAR-assisted VO on two public datasets: [7] and [8]. dataset consists of 27 sequences which are logged with sensors mounted on top of a Segway in driving scenarios. The Segway is equipped with a Ladybug3 panoramic camera, a Velodyne HDL-32E LiDAR, a real-time kinematic (RTK) global positioning system (GPS), etc. Our system takes data from the Velodyne HDL-32E LiDAR and the Ladybug3 panoramic camera as input. For every panoramic image from the camera, we use the distortion parameters provided by [7] to perform distortion correction. The Segway repeatedly explore the campus, both indoors and outdoors, on varying trajectories, at different time and weather. dataset consists of 2 sequences which are logged with sensors mounted on top of a vehicle in driving scenarios. The vehicle is equipped with a Ladybug3 panoramic camera, a Velodyne HDL-64E LiDAR, a Inertial Navigation System(INS) with GPS, etc. Our system takes data from the Velodyne HDL-64E LiDAR and the Ladybug3 panoramic camera as input. For every panoramic image from the camera, we use the distortion parameters provided by [7, 8] to perform distortion correction. We select 6 representative sequences including different time and weather from , and all sequences from for evaluation. The details of all the 8 sequences, indluding time and weather, are listed in Table I. Since the dataset does not explicitly state the time and weather involved in sequences, we do not record them in Table I.
@@ VI-A Comparison with State-of-the-Arts
Table II records the pose estimation results of all testing approaches and our panoramic direct LiDAR-assisted VO. Due to the challenging environment for both VO and LiDAR-assisted VO, no approach is able to run entirely on any of selected sequences. Therefore, we disassemble each sequence into sevral segments and evaluate the testing approaches on these segments. Results in Table II demonstrate that our panoramic direct LiDAR-assisted VO outperforms visual module outperforms all testing approaches on most segments in terms of smaller RMSE of ATE. PAN-SLAM fails to run entirely on 6 segments and achieves large ATE results on 7 sequences. This indicates that even with the full utilization of a 360-degree panoramic image, the performance of VO remains fragile in challenging scenarios if lacking depth sensors to provide distance information. In contrast, our panoramic LiDAR-assisted VO achieves smaller ATE across almost all testing segments, demonstrating the advantage of integrating panoramic camera with 3D LiDAR. On the other hand, LIMO fails to run entirely on 15 sequences and achieves large ATE results on 5 sequeces, while SDV-LOAM fails to run entirely on 10 sequences and achieves large ATE results on 10 sequences. This result indicates that even with the distance information perceived by 3D LiDAR, VO still cannot run stably and robustly if the available information from images are insufficient.
@@ VI-C Impact of Constraints from Different Views
As mentioned in Sec. V-A, constructing phomometric constraints from different views can take full advantage of more available information from panoramic images, and in turn improve the performance of pose estimation. In this section, we design the ablation study of estimating poses with vs. without constraints from different views. Results in Table III demonstrate that utilizing constraints from different views can improve the accuracy of pose estimation on most segments of sequences 2012-11-04. Due to the limitation of page number, we only present the results on one sequence, and the other sequences yielded similar results.
@@ VI-D Visualization
Fig.6 compares the estimated of the trajectories of almost all testing approaches with ground truth on segment 2-2. The reason for not plotting the trajectory of SDV-LOAM is that its estimated trajectory on this segment is too poor to provide any references. In the remaining visualization comparison results, the trajectory estimated by our system is almost entirely overlaid with the ground truth, which further demonstrates the superiority of our approach. When the utilized image is reduced from surround panoramic view to solely forward-facing view, the accuracy of estimated trajectory significantly deteriorates. The location experiencing excessive exposure in the forward-facing view (denoted as lable 1 in Fig. 5) corresponds to the area delineated by the red rectangle in Fig. 6. LIMO and our system utilizing only single-view images, which solely rely on the foreground images for pose estimation, encounters pronounced trajectory disorientation at this situation. In comparison, our system utilizing panoramic-view images and PAN-SLAM maintains a smooth trajectory even in such challenging situations, further demonstrating the significant role of panoramic-view in enhancing the robustness of VO and LiDAR-assisted VO.

### v2

@@ Abstract
Enhancing visual odometry by exploiting sparse depth measurements from LiDAR is a promising solution for improving tracking accuracy of an odometry. Most existing works utilize a monocular pinhole camera, yet could suffer from poor robustness due to less available information from limited field-of-view (FOV). This paper proposes a panoramic direct LiDAR-assisted visual odometry, which fully associates the 360-degree FOV LiDAR points with the 360-degree FOV panoramic image datas. 360-degree FOV panoramic images can provide more available information, which can compensate inaccurate pose estimation caused by insufficient texture or motion blur from a single view. In addition to constraints between a specific view at different times, constraints can also be built between different views at the same moment. Experimental results on public datasets demonstrate the benefit of our panoramic direct LiDAR-assisted visual odometry to state-of-the-art approaches.
@@ I Introduction
Using 3D light detection and ranging (LiDAR) information to assist visual odometry (VO), i.e., LiDAR-assisted VO, has attracted increasing interests recently due to the ability of achieving accurate pose estimation in outdoor environments. Several notable LiDAR-assisted VO systems [1, 2, 3, 4, 5, 6] fuse measurements from a monocular pinhole camera and a 360-degree field-of-view (FOV) LiDAR to perform pose estimation. These approaches suffer from poor robustness if the single view appears motion blur or insufficient texture. Compared to the monocular camera, the panoramic camera provides a 360-degree FOV image data (as shown in Fig. 1), which can more fully associate images with LiDAR points and hold great potential to enhance the robustness of LiDAR-assisted VO. Compared to utilizing rigidly coupled multi-camera systems to achieve 360-degree FOV, the panoramic camera has the advantage of not requiring online extrinsic calibration and temporal synchronization between multiple cameras.
Despite a panoramic camera can provide 360-degree FOV through multiple synchronized surround-view cameras, each camera has a low frame rate and a very limited horizontal FOV. When the camera moves quickly, there is little overlap between consecutive frames from the same perspective. The limited overlap results in less information available for constructing constraints, which in turn leads to poor accuracy in pose estimation.
@@ II Related Work
Wide FOV camera-based VO. Zhang et. al. [9] conducted an in-depth investigation into the impact of camera FOV on the performance of VO, and pointed out that it is advantageous to use a large FOV camera (e.g., fisheye camera) for indoor scenes. Caruso et. al. [10] proposed a direct monocular SLAM method for omnidirectional or wide FOV fisheye cameras, which allow to observe and reconstruct a larger portion of the surrounding environment, and also make the system more robust to degenerate (rotation-only) motion. Matsuki et. al. [11] proposed a direct monocular visual odometry for omnidirectional cameras. [11] utilized the unified omnidirectional model as a projection function, which can be applied to fisheye cameras with a FOV well above 180 degrees. CubemapSLAM [12] presented a real-time feature-based SLAM system for fisheye cameras featured by a large FOV, which increase visual overlap between consecutive frames and capture more pixels belonging to the static parts of the environment. PALVO [13] applied panoramic annular lens to visual odometry, greatly increasing the robustness to rapid motion and dynamic scenarios.
@@
LiDAR-assisted VO. DEMO [1] proposed a feature-based LiDAR-assisted VO, which associates extracted 2D features with 3D points in order to assign depth value to each feature point (i.e., 2D-3D data association). LIMO [2] added a loop closure module on the base of DEMO and utilizes semantic information to identify moving objects and reject outliers. Huang et. al. [3] proposed to use line features in addition to point features for tracking and mapping, so as to improve the robustness to noises [24], large viewpoint changes [25], and motion blurs [26]. DVL-SLAM [4] proposed to use the direct method for LiDAR-assisted visual tracking, aiming to avoid interpolation error in 2D-3D data association.
@@ III-A Coordinate Systems
There are several types of panoramic cameras, and we specifically refer to the type that contains five surround-view cameras in this work (i.e., ). We denote , , as a 3D point in the camera coordinate, the body coordinate and the world coordinate respectively. The body coordinate is coinciding with , and is the transformation form to . The world coordinate system coincides with the body coordinate system at the initial time.
@@ III-B Camera Projection Model
where represents the depth of 2D point in the camera coordinate.
@@ VI Experiments
In this section, we evaluate our panoramic direct LiDAR-assisted VO on two public datasets: [7] and [8]. dataset consists of 27 sequences which are logged with sensors mounted on top of a Segway in driving scenarios. The Segway is equipped with a Ladybug3 panoramic camera, a Velodyne HDL-32E LiDAR, a real-time kinematic (RTK) global positioning system (GPS), etc. Our system takes data from the Velodyne HDL-32E LiDAR and the Ladybug3 panoramic camera as input. For every panoramic image from the camera, we use the distortion parameters provided by [7] to perform distortion correction. The Segway repeatedly explore the campus, both indoors and outdoors, on varying trajectories, at different time and weather. dataset consists of 2 sequences which are logged with sensors mounted on top of a vehicle in driving scenarios. The vehicle is equipped with a Ladybug3 panoramic camera, a Velodyne HDL-64E LiDAR, a Inertial Navigation System(INS) with GPS, etc. Our system takes data from the Velodyne HDL-64E LiDAR and the Ladybug3 panoramic camera as input. For every panoramic image from the camera, we use the distortion parameters provided by [8] to perform distortion correction.
We select 6 representative sequences including different time and weather from , and all sequences from for evaluation. The details of all the 8 sequences, indluding time and weather, are listed in Table I. Since the dataset does not explicitly state the time and weather involved in sequences, we do not record them in Table I.
@@ VI-A Comparison with State-of-the-Arts
Table II records the pose estimation results of all testing approaches and our panoramic direct LiDAR-assisted VO. Due to the challenging environment for both VO and LiDAR-assisted VO, no approach is able to run entirely on any selected sequences. Therefore, we disassemble each sequence into sevral segments and evaluate the testing approaches on these segments. Results in Table II demonstrate that our panoramic direct LiDAR-assisted VO outperforms visual module outperforms all testing approaches on most segments in terms of smaller RMSE of ATE. PAN-SLAM fails to run entirely on 6 segments and achieves large ATE results on 7 sequences. This indicates that even with the full utilization of a 360-degree panoramic image, the performance of VO remains fragile in challenging scenarios if lacking depth sensors to provide distance information. In contrast, our panoramic LiDAR-assisted VO achieves smaller ATE across almost all testing segments, demonstrating the advantage of integrating panoramic camera with 3D LiDAR. On the other hand, LIMO fails to run entirely on 15 sequences and achieves large ATE results on 5 sequeces, while SDV-LOAM fails to run entirely on 10 sequences and achieves large ATE results on 10 sequences. This result indicates that even with the distance information perceived by 3D LiDAR, VO still cannot run stably and robustly if the available information from images are insufficient.
@@ VI-C Impact of Constraints from Different Views
As mentioned in Sec. V-A, constructing phomometric constraints from different views can take full advantage of more available information from panoramic images, and in turn improve the performance of pose estimation. In this section, we design the ablation study of estimating poses with vs. without constraints from different views. Results in Table III demonstrate that utilizing constraints from different views can improve the accuracy of pose estimation on most segments of sequence 2012-11-04. Due to the limitation of page number, we only present the results on one sequence, and the other sequences yielded similar results.
@@ VI-D Visualization
Fig.6 compares the estimated trajectories of almost all testing approaches with ground truth on segment 2-2. The reason for not plotting the trajectory of SDV-LOAM is that its estimated trajectory on this segment is too poor to provide any references. In the remaining visualization comparison results, the trajectory estimated by our system is almost entirely overlaid with the ground truth, which further demonstrates the superiority of our approach. When the utilized image is reduced from surround panoramic view to solely forward-facing view, the accuracy of estimated trajectory significantly deteriorates. The location experiencing excessive exposure in the forward-facing view (denoted as lable 1 in Fig. 5) corresponds to the area delineated by the red rectangle in Fig. 6. LIMO and our system utilizing only single-view images, which solely rely on the foreground images for pose estimation, encounters pronounced trajectory disorientation at this situation. In comparison, our system utilizing panoramic-view images and PAN-SLAM maintain a smooth trajectory even in such challenging situations, further demonstrating the significant role of panoramic-view in enhancing the robustness of VO and LiDAR-assisted VO.